OpenStudy (anonymous):

A question on vectors @ganeshie8

OpenStudy (anonymous):

If A=i+j+k B= -3i +4j -2k Find the angle between this vectors

ganeshie8 (ganeshie8):

\(\large \vec{A} .\vec{ B} = |\vec{A}| |\vec{B}| \cos (\theta)\) seen this formula for dot product before ?

OpenStudy (anonymous):

Is it applicable for three directions

OpenStudy (anonymous):

Well , then how to handle the L.H.S R.H.S is trivial

ganeshie8 (ganeshie8):

thats a good question ! angle and lengths will remain same irrespective of whether you're working them in 2D or 3D

ganeshie8 (ganeshie8):

for LHS we use algebra definition of dot product : \[\large \langle a_1, a_2, a_3\rangle\bullet \langle \color{red}{b_1, b_2, b_3}\rangle = a_1\color{red}{b_1} + a_2\color{red}{b_2} + a_3\color{red}{b_3}\]

ganeshie8 (ganeshie8):

multiply component by component add them

ganeshie8 (ganeshie8):

If A=i+j+k = \(\langle 1, 1, 1\rangle \) B= -3i +4j -2k = \(\langle -3, 4, -2\rangle \)

ganeshie8 (ganeshie8):

\[\large \vec{A} \bullet \vec{ B} = |\vec{A}| |\vec{B}| \cos (\theta)\]

ganeshie8 (ganeshie8):

\[\large \langle 1,1,1 \rangle \bullet \langle -3,4,-2 \rangle = | \langle 1,1,1 \rangle|~ |\langle -3,4,-2 \rangle | \cos (\theta)\]

ganeshie8 (ganeshie8):

evaluate the LHS..

OpenStudy (anonymous):

-1

ganeshie8 (ganeshie8):

yep ! \[\large \langle 1,1,1 \rangle \bullet \langle -3,4,-2 \rangle = | \langle 1,1,1 \rangle|~ |\langle -3,4,-2 \rangle | \cos (\theta)\] \[\large -1 = | \langle 1,1,1 \rangle|~ |\langle -3,4,-2 \rangle | \cos (\theta)\]

ganeshie8 (ganeshie8):

what about the RHS

OpenStudy (anonymous):